OpenStudy (anonymous):

In the trivia bowl ,each finalist must answer four questions correctly. Each question is worth twice as much as the question before it. the fourth is worth $6000. How much is the first question worth? PLEASE HELP ME.

OpenStudy (hoblos):

4th = 1st* 2^3 1st = 4th/8 = 6000/8 = $750